5.) Find the equation for the plane containing the points P (5,1,4), 0 (5.312), R (1,-2,2) --
6.) Build a parametric equations for the line tangent to the curve r(t) = 572; +(6+1); - +3K at the point P (20.-11,8)
Answer #1

P = (5, 1.4) Q = (5, 3.2) R = (1,-2, 2) The position vectors of the points P, Q, R are respectively ă - OP = 5î tŷ +4h 3:00 =ñ -a = (xîty + k) - (5 î+ 7 + A2) = (2-5)i + (y-2)] + (2-4) k F-a) (6-2)x(2-2)] = 0 = (x-5). 2+ (y-1). (-8) +(2-4).8=0 2x -10r(t) = 5 + 2 ît (6t+1)] - t P (20,-11,8) r (t) = lot î+ 69 - 3+²k The point P (20,-11,8) corresponds to t=2 p (2) = 10.2 i